Find two vectors orthogonal to the given vector. u = (-5, -8) Answer How to enter your answer Keypad Keyboard Shortcuts
Answer #1

ų: 2-5,-8) another vector v You-o =) two rector care Gothogonal VE ca, bo (-51-83 ) catib)=0 - 5a-8b=0 a= at -> a= let t=1 ul